T he sum of the ages of two boys is 50 . In 5 years time the elder brother will be four times the younger brother . Find their ages
Answer by mananth(16946)


x+y=50
After 5 years
ages
x+5
y+5
y+5=2(x+5)
y+5=2x+10
y-2x=5
x= one boy's age
y= second boy's age

1 x + 1 y = 50 .............1

-2 x + 1 y = 5 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-1 x -1 y = -50
-2 x + 1 y = 5
Add the two equations
-3 x = -45
/ -3
x = 15
plug value of x in (1)
1 x + 1 y = 50
15 + y = 50
y = 50
y = 35
y = 35
x= 15 one boy's age
y= 35 second boy's age
m.ananth@hotmail.ca